OpenStudy (jacobciezki):

Solve for x. Express the answer in simplest form.1/7=3/4x-1

OpenStudy (mismycatmin):

1/7 = 3/4x - 1 Add 1 to both sides: 1 1/7 = 3/4x Divide 3/4 to both sides or multiply by its reciprocal, 4/3: 1 1/7 * 4/3 = x Convert 1 1/7 to an improper fraction: 8/7 * 4/3 = x Multiply the numerators and denominators together: 32/21 = x or 1/7 = 3-1/4x cross multiply 21-7 = 4x 16 = 4x x= 4